Deal Leasing leased equipment to Hand Company on January 1, 2018. The lease payments were calculated to provide the lessor a 10% return. Nine annual lease payments of $65,000 are due at the beginning of each year beginning January 1, 2018. The present value of an annuity due of $1 at 10 for Nine periods is 6.33493.

Please provide journa entry for Record the lease, Record the cash payment, Record accrued interest, and Record the amortization expense
Answer #1

Jan 1 .Asset Account Dr 411770

               To Lease Liability Account 411770

         (To record the lease 65000 x 6.33493 =411770.45)

Jan 1 Lease Liability Account Dr 65000

               To Cash Account                 65000

          (To Record the cash payment )

Dec 31 Finance Charges Dr 34677

                  To Lease Liability Account 34677

           To record interest (411770 - 65000) x 10 % =34677

Dec 31 Depreciation Account Dr 45752

                           To Asset Account      45752

            To record amortization expense = 411770 / 9 = 45752